OpenStudy (anonymous):

The ordered pair (1, -5) is a solution for the system x + 2y = -9 and x - y = 6. true or false?

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

true

OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5? (2, -1) (0, 5) (5, 0) (1, 7)

OpenStudy (cggurumanjunath):

substitute each of the 4 values say x=2 y=-1 in y=-3x+5 and y=2x+5 if lhs =rhs then it is a solution

OpenStudy (cggurumanjunath):

lhs=left hand side rhs=right hand side

OpenStudy (anonymous):

Using elimination to solve the system 2x + y = 8 and x + y = 4, which of the following could be the first step? x = 4 2x + y = x + y 3x + 2y = 12 2x + (4 - x) = 8 2x + y = 8--------> 2x + y = 8 (x + y = 4)(-1)---> -x - y = -4

OpenStudy (anonymous):

@CGGURUMANJUNATH